• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 11:46:45 on September11,2005

    Entry number 154473

    keyword=controlled access work

    Since approx. 10:30, Dave Meekins, Scot Spiegel and RadCon (Zach)
    have been in Hall A for the target swap out. Vacuum pump down to
    occur overnight with target cooldown Monday DAY.

    No shifts until Monday DAY

    I paged and talked to Sirish (Compton) and Arne Freyberger (cavities).
    If they do something to their systems it should be halog'ed.